A rectangular parking lot has a length that is 12 yards greater than the width. The area of the parking lot is 540 square yards. Find the length and the width. Use the formula, area=Length*width.
Answer by amarjeeth123(569)

Let the width be x.
Then the length is (x+12).
Area=length*width
x(x+12)=540
x^2+12x-540=0
x^2+30x-18x-540=0
x(x+30)-18(x+30)=0
(x-18)(x+30)=0
x=18 is the only positive solution.
The width is 18 yards.
The length is 30 yards.